OpenStudy (anonymous):

a force of 591lb is required to pull a boat up a ramp inclined at 17.0 degreee with the horizontal how much does the boat weight help me plz satellite73 plzzz

OpenStudy (anonymous):

i got 618 for the answer

OpenStudy (anonymous):

i have no idea how to do this did the link not help?

OpenStudy (anonymous):

ya but its showe me that i got 618 i did the same way that they did

OpenStudy (anonymous):

did you draw a free body diagram? and 618 what?

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

i dont no how to

OpenStudy (anonymous):



OpenStudy (anonymous):

ya and how do i solve it now

OpenStudy (anonymous):

sir plz help me i been trying this for 2 hours

OpenStudy (anonymous):

\[Fn=fg(cosx)\] sorry i forgot to include the force of parallel which is the line up the ramp \[F//=fg(sinx)\] that is the more important one. the only unknown value is mass so you know your force of parallel is 591 pounds (i always used newtons). your angle is 17 degrees to get Fg you would M*9.8=Fg. does that help?

OpenStudy (anonymous):

ya but how do i put the equation together

OpenStudy (anonymous):

591=(m(9.8))(sin(17))

OpenStudy (anonymous):

whats m for

OpenStudy (anonymous):

mass. i dont know why your using pounds and weight, but i never have in physics

OpenStudy (anonymous):

you should know that gravity accelerates you perpendicular to the earths surface at 9.8 m/s^2 and mass times acceleration=force.

OpenStudy (anonymous):

591=(m(9.8))(sin(17)) how do i solve this

OpenStudy (anonymous):

591=(m(9.8))(sin(17))

OpenStudy (anonymous):

sir plz

OpenStudy (anonymous):

how do i solve this problem

OpenStudy (anonymous):

i donno why but that link gives an interesting way to come up with the solution. ive never seen this solution in physics. let me check this problem out real fast. because physics you have to use the right units for it to work out

OpenStudy (anonymous):

so the way they solve it is W=591/cos(73)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is your problem?

OpenStudy (anonymous):

a force of 591lb is required to pull a boat up a ramp inclined at 17.0 degreee with the horizontal how much does the boat weight

OpenStudy (anonymous):

i know that. do you not know what cos 73 is?

OpenStudy (anonymous):

.2923717047

OpenStudy (anonymous):

ok now what is 591/.29

OpenStudy (anonymous):

2037.931034

OpenStudy (anonymous):

sounds like you have it to me

OpenStudy (anonymous):

so the answers is 2038

OpenStudy (anonymous):

sir its it right 2038

OpenStudy (anonymous):

according to that page you linked it is and as long as noone cares about sig figs. ive never solved a vector problem like that with such disregard to units before so i couldnt swear to it. i just followed the link you gave me.

OpenStudy (anonymous):

oh k so the final it 2038 thanks you